Find the market equilibrium point for the following demand and supply functions.

Demand:     p = −2q + 290
Supply:     p = 8q + 2

(q, p) =

Demand:     2p = −q + 88
Supply: 3p − q = 72

(q, p) =

Demand:     p = −5q + 220
Supply:     p = 16q + 10
Answer #1

1. p = -2q + 290

    p = 8q + 2

At equilibrium,

Demand = Supply

-2q + 290 = 8q +2

10q = 290 - 2 = 288

q = 288 / 10 = 28.8          [This is equilibrium q]

p = 8q + 2 = 8(28.8) + 2 = $232.4   [This is equilibrium p]

2. 2p = -q + 88

    q = 88 -2p [This is direct demand function]

    3p - q = 72

   q = 3p - 72 [This is direct supply function]

At equilibrium,

Demand = Supply

88 -2p = 3p - 72

5p = 88 + 72 = 160

p = 160 / 5 = $32     [This is equilibrium p]

q = 88 - 2p = 88 - 2(32) = 24   [This is equilibrium q]

3. p = -5q + 220

    p = 16q + 10

At equilibrium,

Demand = Supply

-5q + 220 = 16q + 10

21q = 220 - 10 = 210

q = 210 / 21 = 10          [This is equilibrium q]

p = 16q + 10 = 16(10) + 10 = $170 [This is equilibrium p]